please help. thank you!!The Gilster Company, a machine tooling firm, has several plants. One plant, located in St. Cloud, Minnesota, uses a job orderComplete this question by entering your answers in the tabs below. Reg A1 Reg A2 Assume the St. Cloud plant uses a single plaReg A1 Reg A2 What is the expected cost per unit produced for job no. 110? (Round your intermediate calculations and final anReq B1 Req B2 Req B3 Find the plant wide overhead rate by using expected machine hours. (Round your answer to 2 decimal placeReq B1 Req B2 Reg B3 Find the department overhead rate using expected machine hours for Department A and Department B. (RoundComplete this question by entering your answers in the tabs below. Reg B1 Req B2 Req B3 Calculate the projected manufacturing
Answer #1

Requirement A1:

Plantwide overhead rate = (Departmental overheads + Plantwide overheads) / Total direct labor hours

Departmental overheads = $133,000 + $288,000 = $421,000

Plantwide overheads = $250,000

Total direct labor hours = 46,000 + 15,000 = 61,000

Plantwide overhead rate = ($421,000 + $250,000) / 61,000 direct labor hours = $11.00 per direct labor hour

Requirement A2:

Manufacturing cost per Unit Produced for Job no. 110

Particulars $
Direct Materials 19,000
Direct labor cost ($33,000 + $7,600) 40,600
Overhead (2,200 + 1,200) * $11.00 37,400
Total Manufacturing cost 97,000

Manufacturing cost per unit = $97,000 / 15,000 units = $6.47 per unit

Requirement B1:

Plantwide overhead rate = Plantwide overheads) / Total machine hours

Plantwide overheads = $250,000

Total machine hours = 19,000 + 48,000 = 67,000 machine hours

Plantwide overhead rate = $250,000 / 67,000 direct labor hours = $3.73 per machine hour

Requirement B2:

Departmental overhead rate = Departmental overhead cost / Departmental machine hours

Department A:

Departmental overhead rate = $133,000 / 19,000 machine hours = $7.00 per machine hour

Department B:

Departmental overhead rate = $288,000 / 48,000 machine hours = $6.00 per machine hour

Requirement B3:

Particulars $
Direct Materials 19,000
Direct labor cost ($33,000 + $7,600) 40,600
Plantwide Overhead (120 + 1,200) * $3.73 4,924
Departmental Overhead (120 * $7.00) + (1,200 * $6.00) 8,040
Total Manufacturing cost 72,564

Manufacturing cost per unit = $72,564 / 15,000 units = $4.84 per unit

All the best...